The second hand of a clock is 1m long . Its angulat velociry is

Grade:11

1 Answers

Ankit Yadav
6 years ago
The hour hand of clock rotates full circle in 12 hours. So it covers an angle 360⁰ in 12 hours. since hour hand of clock covers an angle of 360° in 12hours,Therefore second hand of clock will cover a full circle in 12hrs,=2πThe length of hand has no effect on angular velocity coz it is expressed in terms of radians/sec....2π/60=π/30=6°/sec.Hope you get it